### v1

# Towards Unifying Interpretability and Control: Evaluation via Intervention

Abstract
With the growing complexity and capability of large language models (LLMs), a need to understand model reasoning has emerged, often motivated by an underlying goal of controlling and aligning models. While numerous interpretability and steering methods have been proposed as solutions, they are typically designed either for understanding or for control, seldom addressing both, with the connection between interpretation and control more broadly remaining tenuous. Additionally, the lack of standardized applications, motivations, and evaluation metrics makes it difficult to assess these methods’ practical utility and efficacy. To address the aforementioned issues, we propose intervention as a fundamental goal of interpretability and introduce success criteria to evaluate how well methods are able to control model behavior through interventions. We unify and extend four popular interpretability methods—sparse autoencoders, logit lens, tuned lens, and probing—into an abstract encoder-decoder framework. This framework maps intermediate latent representations to human-interpretable feature spaces, enabling interventions on these interpretable features, which can then be mapped back to latent representations to control model outputs. We introduce two new evaluation metrics: intervention success rate and the coherence-intervention tradeoff, designed to measure the accuracy of explanations and their utility in controlling model behavior. Our findings reveal that (1) although current methods allow for intervention, they are inconsistent across various models and features, (2) lens-based methods outperform others in achieving simple, concrete interventions, and (3) interventions often compromise model performance and coherence, underperforming simpler alternatives, such as prompting, for steering model behavior and highlighting a critical shortcoming of current interpretability approaches in real-world applications requiring control. Code is made available for replicability.
1 Introduction
As large language models (LLMs) have become more capable and complex, there has emerged a need to better understand and control these models to ensure their outputs are safe and human-aligned. Many interpretability methods aim to address this problem by analyzing model representations, attempting to understand their underlying computational and reasoning processes in order to ultimately control model behaviour. While many of these methods, and interpretability as a field more broadly, claim control and intervention as abstract goals and present compelling qualitative results demonstrating that intervention is possible (for example, Anthropic’s Golden Gate Claude Anthropic ; Templeton et al. (2024)), the link between interpretation and intervention is often tenuous in practice, as many methods are not explicitly tailored for both. Furthermore, even fewer are thoroughly and systematically evaluated for the ability to control model outputs beyond simple qualitative examples. We believe the reason for this is threefold. First, interpretability methods produce explanations in disparate feature spaces, such as the token vocabulary, probe predictions, or learned auto-interpreted features, hindering comparisons across methods. Second, there exists a “predict/control discrepancy” (Wattenberg & Viégas, 2024), where the features identified by interpretability methods for predicting behavior are not the same as those used for steering it. Third, there do not exist standard systematic benchmarks to measure intervention success.
Figure 1: Our proposed intervention framework, which encodes model latent representations, $x$, into human-interpretable features, $z=xD$, that can then be perturbed to $z^{\prime}$ and decoded back into counterfactual latent representations, $\hat{x}^{\prime}$.
In this work, we view intervention as a fundamental goal of interpretability, and propose to measure the correctness of interpretability methods by their ability to successfully edit model behaviour. In particular, we focus on sparse autoencoders  (Cunningham et al., 2023; Rajamanoharan et al., 2024a; Templeton et al., 2024; Bricken et al., 2023; Gao et al., 2024), Logit Lens  (nostalgebraist, ; Dar et al., 2023), Tuned Lens  (Belrose et al., 2023; Din et al., 2023), and linear probing  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022), and benchmark them with steering vectors and prompting as baselines for intervention. In order to enable comparison across these various methods, we first unify and extend the methods as instances of an abstract encoder-decoder framework, where each method encodes uninterpretable latent representations of language models into human-interpretable features and the decoder of the framework inverts this mapping, allowing us to reconstruct a latent representation from the features. Under this abstract framework, we can intervene on the interpretable feature activations generated by each method and decode them into latent counterfactuals, which produce counterfactual outputs corresponding to the desired intervention.
The unifying feature interpretation and intervention framework allows us to propose two standard metrics for evaluating mechanistic interpretability methods: (1) intervention success rate, which measures how well intervening on an interpretable feature causally results in the desired behavior in the model outputs, and (2) coherence-intervention tradeoff, which measures how well the causal interventions succeed without damaging the coherence of the model’s outputs. We evaluate Logit Lens, Tuned Lens, sparse autoencoders, and linear probes, for these metrics on GPT2-small, Gemma2-2b, and Llama2-7b, comparing them to simpler but uninterpretable baselines of steering vectors and prompting. Our results show that while existing methods allow for intervention, they are inconsistent across features and models. Furthermore, lens-based methods outperform all other methods including sparse autoencoders for simple, concrete features. This is likely due to frequent mislabeling of the learned SAE features, where a feature labeled “references to coffee” actually seems to encode for ‘beans’ and ‘coffee beans’ more than ‘coffee’. Similarly, probes and steering vectors are prone to learning spurious correlations in their training data, whereas Logit Lens and Tuned lens do not suffer from the same sensitivities. Finally, we show that intervention often comes at the cost of model output coherence, underperforming simple prompting baselines, presenting a critical shortcoming of existing methods in real-world applications that require control and intervention. We conclude this work with detailed takeaways about the strengths and weaknesses of each method, including discussion of which methods are optimal for specific intervention topics, which are best to use out of the box, and which hold the most promise for future development.
Our main contributions include:
  • In Section 3.1, we present an encoder-decoder framework that unifies four popular mechanistic interpretability methods: sparse autoencoders, logit lens, tuned lens, and probing.
  • In Section 3.2, we propose two evaluation metrics for encoder-decoder interpretability methods, namely (1) intervention success rate; and (2) the coherence-intervention tradeoff to evaluate the ability of interpretability methods to control model behavior. We also provide a dataset for benchmarking state-of-the-art LLMs on these metrics.
  • In Section 4, we perform experimental analysis on GPT-2, Gemma2-2b, and LLama2-7b, and present detailed takeaways comparing interpretability- and control-based methods.
We hope that this paper encourages future work in the interpretability community to make explicit the intent and application of methods, particularly in reference to control and intervention, and to evaluate and benchmark causality and utility in application.
2 Related Work
Mechanistic Interpretability. Existing work in mechanistic interpretability broadly falls into two categories: activation patching and interpreting hidden representations. Activation patching utilizes carefully constructed counterfactual representations to study which neurons or activations play key roles in model computation, ideally localizing specific information to individual layers, token positions, and paths in the model (Geiger et al., 2021; Vig et al., 2020). However, recent work points to key limitations of patching, particularly with respect to real-world utility in downstream applications such as model editing (Hase et al., 2024; Zhang & Nanda, 2023). As such, we focus primarily on methods for inspecting hidden representations, of which probes are the most commonly used  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022). Other methods such as nostalgebraist ; Dar et al. (2023) project intermediate representation into the token vocabulary space, with Belrose et al. (2023); Din et al. (2023); Geva et al. (2022) building and improving upon these early-decoding strategies. Ghandeharioun et al. (2024a) unifies most of these methods into an abstracted framework for inspecting model computation. More recently, sparse autoencoders and dictionary learning have been explored as a solution to the uninterpretability of model neurons, particularly due to issues with polysemanticity and superposition (Elhage et al., ; Bricken et al., 2023; Cunningham et al., 2023; Bhalla et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024a; Templeton et al., 2024; Karvonen et al., 2024; Dunefsky et al., 2024).
Evaluation. Due to the recency of the field, standard evaluation metrics across interpretability methods have not yet been established, and similar to the broader interpretability field, evaluation is frequently ad-hoc and primarily qualitative in nature. With regards to quantitative metrics, in Gao et al. (2024); Templeton et al. (2024); Makelov et al. (2024), sparse autoencoders are evaluated for reconstruction error, recovery of supervised or known features, activation precision, and the effects of ablation; however, none of these metrics measure the human interpretability or correctness of explanations. Sparse autoencoder evaluation has also sometimes included ratings by humans to determine the interpretability of the learned features for labeling, as done in Rajamanoharan et al. (2024b), but such measures do not necessarily reflect downstream utility in model control.
Causal Intervention. Previous literature on probing frequently evaluates learned probes and features through intervention to ensure causality and correctness, as done in Li et al. (2022); Chen et al. (2024); Hernandez et al. (2023b; a); Marks & Tegmark (2023). The interventions performed for measuring causality are similar to those used to perform model “steering” Rimsky et al. (2023); Panickssery et al. (2024); Ghandeharioun et al. (2024b) and should ideally produce the same effect but with the added claim of interpretability. Beyond probing, only Belrose et al. (2023); Chan et al. (2022); Olah et al. (2020); Templeton et al. (2024) consider causal intervention as a tool for evaluation. However, Belrose et al. (2023); Olah et al. (2020); Chan et al. (2022) mainly focus on mean ablation for intervention, with no exploration of the coherence of the counterfactual outputs or their utility in application. Templeton et al. (2024) on the other hand only provides a qualitative evaluation of intervention via their ‘Golden Gate Claude’ [Anthropic ] and does not compare to other interpretability methods.
3 Method
In this section, we first introduce a unifying framework for four common mechanistic interpretability methods: sparse autoencoders, Logit Lens, Tuned Lens, and probing, along with modifications to these methods that permit principled intervention on representations. We then propose evaluation metrics for (1) testing the correctness of explanations via intervention and (2) the usefulness of these methods for steering and editing representations and model outputs.
3.1 Unifying Intervention Framework
Latent vectors to interpretable features.
The central aspect of most interpretability work is the ability to translate model computation into human-interpretable features, whether the computation be latent directions, neurons, components, reasoning processes, etc. Many works aiming to explain LLMs focus particularly on hidden representations, where the mapping between high-dimensional dense embeddings and human-interpretable features is modeled through a (mostly) linear dictionary projection or affine function:
$z$ $=f(x)=\sigma(x\cdot D)$
$\hat{x}$ $=g(z)\approx f^{-1}(z)=z\cdot D^{-1}$
$z^{\prime}$ $=\text{Edit}(z),~{}~{}\hat{x^{\prime}}=g(z^{\prime})$
where each $z_{i}$ is a feature activation, each $i$ in $D$ corresponds to a human-interpretable feature [^1] While “features” are not necessarily well-defined in the interpretability community, we find that the second definition in Elhage et al. most closely aligns with our own notion of features as human-interpretable concepts that are relevant for manipulation and editing., and $\sigma$ is an activation function that is frequently the identity. In the case of sparse autoencoders, $D$ is a learned, overcomplete dictionary, with $N$ in range 16k - 65k for small models (up to 16M for large models), and $\sigma$ is a ReLU, JumpReLU, or ReLU + top-k activation function. Given that SAE features are learned, they are not immediately interpretable and must be labelled by humans or strong LLMs. For Logit Lens, $D$ is simply the language model’s unembedding matrix, meaning each feature corresponds to a single token in the vocabulary. For Tuned Lens, $D$ is the exact same as Logit Lens but with a learned linear transformation applied. Linear probes can be thought of as a learned dictionary with $N=1$ where $\sigma$ is a sigmoid or softmax activation. Of all these methods, Logit Lens is the only method that does not require any training data, and sparse autoencoders are the only method that does not produce immediately interpretable features. Please see Figure 1 for a visual summary of these methods.
Interpretable features to counterfactual latent vectors.
While producing explanations is straightforward for each of these methods, intervening on model representations using the information provided by explanations is not as simple. Doing so requires defining a reverse mapping from the explanations to the latent representations of the model, which is only explicitly done by sparse autoencoders. We defined the inverse mappings for the rest of the functions as follows. To map Logit Lens’s explanations back into the model’s latent space, we would ideally apply the inverse of the unembedding matrix to $z$; however, in practice this is often ill-conditioned due to the dimensionality of $D$. As such, we instead use the low-rank pseudoinverse of the unembedding matrix and right-multiply it to the explanation logits. Similarly, for Tuned Lens, we model the decoding process through the pseudo-inverse of the Tuned Lens projection applied to the unembedding matrix. Notably, both of these methods only require a simple linear transformation to go back-and-forth between latent vectors and explanations. For probing, $D^{-1}$ can be defined such that $\hat{x}^{\prime}=x+\theta$, as proposed by Chen et al. (2024) which adds the linear probe’s weights to the latent representation, bypassing the need for optimization. Sparse autoencoders have a well-defined backwards mapping through the SAE decoder, which is frequently linear in practice and often the transpose of the encoder weights. Thus, we are now able to decode $z$ into $\hat{x}$ for all methods.
Intervening on interpretable features.
Given the above framework, intervention is performed by directly altering the feature activation $z_{i}$ corresponding to the desired feature $i$ to be edited. While the edited activation $z_{i}^{\prime}$ can naively be set to some constant value $\alpha$, the same constant may have drastically varying effects for different tokens and different prompts. As such, to take into account the context of $z$, for Logit Lens, Tuned Lens, and SAEs we set $z_{i}^{\prime}=\alpha*\max(z)$. This ensures that the feature $i$ is the most dominant feature in the latent vector for $\alpha>1$. Decoding $z^{\prime}$ yields the altered latent representation $\hat{x}^{\prime}=g(z^{\prime})$, which accounts for both the error of the explanation method as well as the intervention performed. For probing and steering vectors, $\hat{x}^{\prime}=x+\alpha*v$, where $v$ is the steering vector or the weights of the linear probe. Note that $\alpha$ is a hyperparameter that must be tuned for each method, model, and sometimes even intervention feature and thus cannot be used to compare the effects of interventions across methods. In order to do so, we can instead measure the normalized difference between the latent vectors $x$ and $\hat{x}^{\prime}$, to better understand the relationship between the degree of intervention performed and our various evaluation metrics. We also note that $\hat{x}$ and $\hat{x}^{\prime}$ are not necessarily in-distribution for the language model, but due to the additive nature of the residual stream, the linear representation hypothesis, and the linear nature of the decoder/unembedding matrix, we believe that such interventions may still be principled in practice (see Park et al. (2023) for more on the linear representation hypothesis and intervention).
Open-ended generation.
In order to generate open-ended text after intervening on the explanation, we edit the corresponding representations in place, as is common practice with prior steering methods. Formally, the representation $x_{t}$ at token position $t$ and layer $l$ is edited to be $\hat{x_{t}}^{\prime}$, ensuring a causal effect on all ensuing tokens $x_{t+1},x_{t+2},...,x_{T}$.
3.2 Comparative Evaluation Across Methods and Models
Given the overall lack of standardized evaluation of mechanistic interpretability methods, we intend for this work to serve as a starting point for systematic evaluation by evaluating methods in simple, easy-to-measure contexts. In particular, we think of our evaluations as measuring a kind of upper bound for these methods: in the easiest of settings, how well do existing methods work?
Explanation Correctness
We first propose metrics to evaluate the correctness of explanations and interventions. More specifically, to test whether a single feature of an explanation $z_{i}$ is correct, we intervene on that feature to produce $z_{i}^{\prime}$ and decode $z^{\prime}$ to $\hat{x}^{\prime}$. We can then hypothesize about the corresponding change in the model’s output when generating from $\hat{x}^{\prime}$ beforehand, as it should match the intervention made to produce $z^{\prime}$. For example, if feature $i$ encodes the concept “references to Paris,” increasing the value of $z_{i}$ should result in increases to references of Paris in the model’s output. From this, we propose a metric of Intervention Success Rate, which measures if increasing the activation of feature $i$ by increasing $z_{i}$ results in the appropriate increase of the feature $i$ in the model’s output. To evaluate a continuous relaxation of this, we can also similarly measure the probability assigned to tokens relating to feature $i$. As such, even if the model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of $i$ results in any change to the model’s output at all. We refer to this metric as Intervened Token Probability. Importantly, both of these metrics can be thought of as measuring the causal fidelity of the features highlighted by explanations.
Usefulness of Intervention Methods
[…]
An Open-ended Evaluation Dataset.
[…]
4 Experiments
In this section, we evaluate the five methods on Gemma2-2b Team et al. (2024), Llama2-7b Touvron et al. (2023), and GPT2-small Radford et al. (2019). We first measure the completeness of the explanations produced in Section 4.2. We then benchmark the methods on the metrics described in Section 3.2, providing quantitative results in Sections 4.3 and 4.4 and qualitative results in Section 5. Finally, we present an analysis of the empirical directional similarities between methods in Section 4.5, and analyses of the efficacy of methods at different model depths in Appendix A.6.
4.1 Implementation Details
Intervention Topics. We pick 10 intervention topics that all relate to references to specific words or phrases: {‘beauty’, ‘chess’, ‘coffee’, ‘dogs’, ‘football’, ‘New York’, ‘pink’, ‘San Francisco’, ‘snow’, ‘yoga’}. These simple, low-level features are ideal for evaluation through intervention for four key reasons: first, measuring the presence of a word or phrase is much easier than measuring a high-level abstract concept such as sycophancy, second, these features were available across many layers for the Gemma2-2b and GPT2-small sparse autoencoders, third, the features necessarily exist in the Logit Lens unembedding dictionary, and finally, datasets that are labelled for the presence of these features are very straightforward to collect for generating steering vectors and probes. As such, we can easily compare intervention on these features across methods and also easily measure intervention success by checking if the given word/phrase exists in the model’s output.
Table 1: Normalized latent reconstruction error without intervention.
Method Gemma2-2b Llama2-7b GPT2-small
Logit Lens 0.52 $5e^{-5}$ 0.22
Tuned Lens $5e^{-3}$ 0.32
SAEs 0.38 1.64
Steering vectors and probing. We implement steering vectors as described in Rimsky et al. (2023) with a few simple modifications. Where in Rimsky et al. (2023), the difference between contrastive pairs is taken only at the last token, we find that averaging across the token dimension and taking the difference between those averages yields much better results. This is due to the fact that in Rimsky et al. (2023), the only difference between representations occurred in the token position of the answer letter, or the last token; however, in our case the information related to the intervention feature could be present at any token. Contrastive data pairs were generated by prompting ChatGPT for 200 pairs sentences or phrases about the desired intervention topic and about other topics within the same genre. For example, we used the prompts “Can you please generate 100 short sentences or phrases with the color ‘pink’ in them?” to create positive samples and “Can you also generate 100 phrases or sentences with other colors (not pink) in them?” to create negative ones. We also use these contrastive pairs to train the linear probes, using the implementation from Chen et al. (2024). All probes reached train and test accuracies of 100% across all models and intervention topics.
Sparse autoencoders. We focus specifically on sparse autoencoders trained to interpret the residual stream of transformer models. We use the SAELens library from Bloom (2024b) for GPT2-small and the Gemma Scope SAEs released by Google in Lieberum et al. (2024) for Gemma2-2b. For both models, SAE feature labels were found via Neuronpedia Lin & Bloom (2023), which allows users to interact with, explore, and search for auto-interpretation labelled learned features. Note that the Gemma SAEs are of the JumpReLU architecture from Rajamanoharan et al. (2024b) whereas the GPT2-small SAEs are standard ReLU + L1 regularization SAEs, as noted in Bloom (2024a).
4.2 Sanity Checking Explanation Reconstructions
Before testing these methods for their ability to intervene, we first want to evaluate the completeness of the explanations and the effect of replacing $x$ with $\hat{x}$ without any intervention or editing. We do so by measuring the normalized latent reconstruction error: $\texttt{Error}=||\hat{x}-x||/||x||$ where $\hat{x}=g(f(x))=g(z)$. This error is a key part of the loss function that sparse autoencoders are trained on and measures the information loss incurred by mapping between the language model’s latent space and the interpretable feature space. Given that steering vectors and linear probes do not output complete explanations, we only measure this error for the other three methods, as shown in Table 1, where we see that errors vary a lot across models but most methods are relatively consistent in their error, with the exception of the GPT2-small sparse autoencoders.
We also measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7. We find that the coherence of the outputs generated by the reconstructed latents generally matches the coherence of the clean model outputs. We use a deviation of ±$\pm 1$ around the mean of clean output coherence scores as a threshold for future evaluations, shown in the dashed lines.
Figure 2: Evaluation of the Intervention Success Rate for each method. Note that normalized edit distance is a proxy for intervention intensity that is comparable across methods. Logit Lens outperforms all other methods across all models.
4.3 Intervention Success Across Models
As described in Section 3.2, in order to evaluate the correctness of explanations, we measure the causal effects of intervening on specific features of each explanation. For a given feature or intervention topic $i$, we see if increasing the activation of that feature results in an increase of the feature in the model’s output in the top row of Figure 2. In order to compare across methods, which all have different explanation feature spaces and scales, we measure the success of interventions as a function of the norm of the distance between the edited latent representation $\hat{x}^{\prime}=g(Edit(f(x)))$ and the original latent representation $x$: $||\hat{x}^{\prime}-x||/||x||$ as done in Section 4.2. Results for intervened token probability are in Appendix A.4.
We find that across all three models, Logit lens and Tuned lens have the highest intervention success rate, regardless of the normalized edit distance. While increasing intervention strength and editing the latent representations more for steering vectors and probes results in greater intervention success for GPT2-small and LLama2-7b, the same is not true for Gemma2-2b, for which intervening with steering vectors and probes seems to have minimal effect. Furthermore, we find that intervening with sparse autoencoders is moderately successful, but due to the large reconstruction error for GPT2-small, even with intervention, the edit distance is much greater than that needed for Logit lens and Tuned lens. In general, we believe the lower performance of Sparse Autoencoders is due to heavy noise in the labels of features. For example, a feature labelled ‘references to coffee’, is sometimes actually a feature that encodes for references to ‘beans’ and ‘coffee beans’, and thus only sometimes increases mentions of ‘coffee’. Probes and steering vectors also have suboptimal performance, often due to learning of spurious correlations in the training data rather than the true intervention feature.
When measuring token probability rather than output presence, as explored in Appendix A.4, we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed. We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
4.4 Effects of Intervention on Output Quality
We next measure the coherence of the intervened output text produced by each method to ensure that intervention through mechanistic interpretability methods is possible in applications and does not damage the utility of the model. We measure coherence as described in Section 3 as a function of the normalized latent edit distance $||\hat{x}^{\prime}-x||/||x||$ and with respect to the intervention success rate in Figure 4. We visualize the mean of coherence scores for the clean model outputs with solid black horizontal lines, the same as those shown in Figure 7, with a buffer of ±$\pm 1$ around the mean in dashed lines. We also consider a prompting baseline, where we simply prompt the language model to constantly talk about the intervention topic, to better understand the optimal coherence possible while satisfying the intervention. This is shown by the teal stars in 4. Given that only Llama2-7b was instruction-tuned out of the models evaluated, prompting was less successful for Gemma2-2b and infeasible for GPT2-small. Also note that the intervention success rate approaches near 100% with prompting as the number of generated tokens increases; however, seeing as we only generate 30 tokens, the success rate is slightly lower than expected.
We find that for GPT2-small and Llama2-7b, intervention with most methods results in a sharp decline in text coherence, with Logit Lens and Tuned Lens having the most dramatic fall-off. In fact, we find that even with a full point drop in coherence from the baseline models’ coherence (denoted with the black lines), the representations can barely be edited by any methods except for Gemma2-2b. While at first, this seems quite concerning, as this would prevent the real-world application of these methods to scenarios necessitating high model performance, viewing coherence as a function of intervention success rate tells a different story, as shown in Figure 4. In fact, we see that Logit Lens and Tuned Lens have relatively good coherence-intervention tradeoffs for GPT2-small and LLama2-7b, with success rates of around 0.5 and 0.6 respectively when considering outputs within one point of deviation from the mean coherence score of the clean model. All other methods exhibit much less desirable Pareto curves. We also note that as the models increase in size and performance, with GPT2-small being the smallest and worst language model and Llama2-7b being the largest and most performant, the Pareto curves are shifted up and to the right more for most methods, suggesting that targeted intervention is more possible as models become more capable.
Figure 3: Intervened model output coherence measured with respect to normalized edit distance. The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.
Figure 4: Intervened output coherence measured with respect to intervention success rate. The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.
Qualitative Examples. In addition to our empirical results, we present examples of the output text for all intervention methods in Figure 5 for qualitative evaluation of intervention on the feature ‘yoga’. Examples for the “Optimal intervention strength” (left column) were randomly chosen from the outputs where intervention succeeded and coherence was still relatively high. Examples for “Excessive Intervention” were randomly chosen from the outputs of the highest intervention strength tested (right column). Note that intervention results in infinite repetition at very high intervention strengths for all methods; however, only Logit Lens and Tuned Lens result in repetition of tokens related to ‘yoga’. More examples are provided in Appendix A.7.
Figure 5: Examples of intervened model outputs for intervention feature ‘yoga’ at both the optimal intervention strength (left) and the maximum intervention strength tested (left). Outputs degrade into incoherent repetition at high intervention strength for all methods.
4.5 Intervention Similarity Between Methods
We finally evaluate the empirical similarity between each of these methods. By converting each intervention into one in the latent space of the model, we can measure the cosine similarity between edit directions for each method. More specifically, for each method and intervention topic, we compute the expectation of $\hat{x}^{\prime}-x$ over all samples. This difference vector can be thought of as the direction of the intervention topic in the model’s latent space. We compute the cosine similarity between these vectors for each intervention topic and display the average across topics in Figure 6.
We find that Logit Lens and Tuned Lens are highly similar, as expected. Similarly, steering vectors and probe weights tend to lie in similar directions, likely due to the same underlying data used to train both. Most interestingly, we find that sparse autoencoders tend to intervene in somewhat similar directions to steering vectors and probes. Even more surprisingly, they have near orthogonal directions to Logit Lens and Tuned Lens, even when interventions succeed for all methods. This is likely due to the high-dimensionality of model latent spaces, allowing for many vectors to be near-orthogonal. We speculate that sparse autoencoders may be more similar to probes and steering vectors because the three methods may have a bias toward representing past information and tokens, due to their training and labelling algorithms. Logit Lens and Tuned lens, on the other hand, are designed to reveal information about the next token specifically, given that they are early-decoding strategies and thus may contain more information about model outputs rather than inputs.
Figure 6: Cosine similarity of intervention directions in model latent space across methods. Logit Lens and Tuned Lens are highly similar, as well as steering vectors and probes.
4.6 High-Level Comparison of Methods
Given the differences between methods, including how their features are learnt, what explanations they produce, and the causal fidelity and utility of explanations for steering, we present the following characterization of the strengths and weaknesses of each method.
Logit Lens is easy to use, requires no training, and maps features directly to vocabulary tokens, making it highly interpretable, and we find that it generally has high causal fidelity as well. However, its predefined, static features are limited and rudimentary, preventing its utility for steering in many real-world or safety-critical applications where abstract features such as truthfulness and bias are more relevant.Tuned Lens shares these traits but requires an additional learned affine transformation, although many are already open-sourced. Sparse autoencoders learn a very large dictionary, allowing for features to be both low-level and high-level or abstract, potentially including safety-critical features, and ideally “covering” all the true underlying features learned by the model. Furthermore, explanations are sparse, increasing their human interpretability. However, the learned features are frequently not human-interpretable, leading to labeling errors and lower causal fidelity in practice. Furthermore, they require significant training and labeling data and compute, and there is no guarantee that a desired feature will exist in the SAE’s dictionary, also preventing their utility in safety-critical domains. Probing allows for feature specification, given that a labeled dataset is available, but is prone to learning spurious correlations, leading to low causal fidelity and minimal steering utility. Steering vectors are very similar to probes in practice, but they require no optimization and are slightly less sensitive to correlations in the training data. As such, we believe that lens-based methods are most useful for providing high-fidelity explanations but are likely not effective solutions for steering in real applications, for which steering vectors and prompting are more promising but require careful oversight and refinement to ensure efficacy. SAEs offer a balance between interpretation- and control-based methods but also suffer from a union of their problems.
5 Conclusion
As language models are integrated into everyday aspects of life, whether benign or life-changing, being able to understand them and ensure they behave as expected is of utmost importance. While interpretability methods show great promise, the correctness of their interpretations is less clear. In this work, we emphasize that intervention is a fundamental goal of interpretability, both to evaluate the correctness of interpretability methods as well as their utility in real-world applications. By unifying and extending four popular interpretability methods — sparse autoencoders, logit lens, tuned lens, and probing — into an abstracted encoder-decoder framework, we enable structured interventions on human-interpretable features, mapfping them back to model latent representations to directly control outputs. Through the introduction of two standardized metrics, the intervention success rate, and the coherence-intervention tradeoff, we provide a rigorous basis for evaluating the causal fidelity and utility of interpretability methods. Our findings reveal significant inconsistencies in intervention capability across different methods and models, with simpler lens-based approaches generally outperforming the more compute-intensive sparse autoencoders for simple, concrete interventions. Furthermore, non-interpretability-based approaches, such as prompting, perform best overall, demonstrating a significant gap in current interpretability approaches.
Acknowledgments
The authors would like to thank Fred Zhang and Lucas Dixon for their helpful discussion and comments on the manuscript. This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, AI2050 Early Career Fellowship by Schmidt Sciences, and faculty research awards from Google, OpenAI, Adobe, JPMorgan, Harvard Data Science Initiative, and the Digital, Data, and Design (Dˆ3) Institute at Harvard. UB is funded by the Kempner Institute Graduate Research Fellowship. The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.
Appendix A Appendix
A.1 Additional Implementation Details: Intervention Hyperparameter $\alpha$
When intervening on $z$ to get $z^{\prime}$ with Logit Lens, Tuned Lens, and SAEs, we set $z_{i}^{\prime}=\alpha*max(z)$. For probing and steering vectors, $\hat{x}^{\prime}=x+\alpha*v$, where $v$ is the steering vector or the weights of the linear probe. Note that $\alpha$ is a hyperparameter that must be tuned for each method and model, and thus cannot be used to compare the effects of interventions across methods. We record the values of $\alpha$ used in our experiments in Table 2.
Table 2: .
Method
GPT2-small
Layer 9
Gemma2-2b
Layer 20
Llama2-7b
Layer 18
Logit Lens [50, 70, 90, 110, 130] [100, 130, 160, 200, 230] [0.5, 3, 7, 11, 15, 19]
Tuned Lens [20, 25, 30, 35, 40] [1, 7, 11, 15, 19, 23]
SAEs [3, 4, 5, 6] [1, 2, 3, 4, 5]
Probing [150, 200, 250, 300, 350] [200, 250, 300, 350] [10, 90, 110, 130, 150]
Steering Vectors [2, 4, 6, 8, 10] [3, 4, 5, 6] [0.5, 3, 4, 5, 6]
A.2 Additional Implementation Details: SAE Features
As described in Section 4.1, we use the sparse autoencoders hosted on SAELens and find the relevant features with Neuronpedia’s exploration and search tools. We document all of the features we consider for each intervention topic in Table 3. Note that for some specified intervention topics, an exact feature match does not exist for the GPT or Gemma SAEs. As such, we either exclude that topic or consider the closest-related topic (such as “instruction related to yoga poses and their benefits” when what we would like is “references to yoga”). Many of these imperfect features still yield reasonable intervention success rates.
Table 3: .
Intervention
Feature
GPT2-small
Layer 9
Feature
GPT2-small SAE
Layer 9 Name
Gemma2-2b
Layer 20
Feature
Gemma2-2b SAE
Layer 20 Feature Label
San Francisco 11233
“mentions of the city
of San Francisco”
3124
“references to San Francisco
and related locations”
New York 5831
“references to the city
of New York”
3761
“specific place names and
geographical locations in New York”
beauty 1805
“words related to beauty
or aesthetic appreciation”
485
“instances of the word
“beauty” in various contexts”
football 11252
“references to football
and baseball contexts”
pink 2415
“mentions of the
word “Pink.””
13703
“references to the color
pink and its various associations”
dogs 12435
“mentions of dogs or
dog-related terms”
12082
“references to dog behavior
and interactions”
yoga 6310
“instructions related to
yoga poses and their benefits”
chess 21685
“mentions of the
game of chess”
13419
“elements within
the context of chess”
snow 5053
“references to
snow-related terms”
13267
“references to snow and
related terms”
coffee 23472
“references to
coffee-related words”
15907
“references to coffee and
related cafés or establishments”
A.3 Additional Evaluations: Coherence of Method Outputs without Intervention
We measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7, which we can compare to the baseline of the clean model outputs (labelled ‘Clean’ and shown in black). We find that the coherence of the outputs generated by the reconstructed latents generally matches the coherence of the clean model outputs. We use a deviation of ±$\pm 1$ around the mean of clean output coherence scores as a threshold for future evaluations, shown in the dashed lines.
Figure 7: Histogram of coherence scores for clean model outputs (Clean) and for the models where $x$ is replaced by $\hat{x}$ without any intervention for Logit Lens, Tuned Lens, and SAEs. Dashed lines show the mean for each distribution.
A.4 Additional Metrics: Intervened Token Probability
Please see Section 3.2 for more details. We measure the probability assigned to tokens relating to feature $i$ when intervening on feature $i$. As such, even if a model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of feature $i$ results in any change to the model’s output at all. We refer to this metric as Intervened Token Probability. Importantly, both of these metrics can be thought of as measuring the causality of the features highlighted by explanations.
Results for Intervened Token Probability are shown in 8, where we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed. We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
Figure 8:
A.5 Additional Metrics: Perplexity
As described in Section 3.2, we evaluate the perplexity of the intervened generated text to measure the utility of interpretability methods for targeted intervention in 9. We measure this perplexity with respect to a stronger language model than the one studied, in this case with Llama3.1-8b.
We find that the results for perplexity are generally unintuitive and do not align with the results for coherence. We hypothesize that perplexity is not a useful measure when text is extremely out-of-distribution with respect to normal text, and in particular when the text is highly repetitive. For example, if the same token is repeated 20 times, we (and other language models) might assume that the next 20 tokens would also be the same, resulting in a low perplexity even if the quality of the text is poor. As such, we do not consider these results to be particularly meaningful or significant.
Figure 9:
A.6 Intervention Efficacy Across Model Depth
In order to ensure the generalizability of the above results across layer depths, we repeat all experiments for each layer of GPT2-small, as shown in FIgure 10. However, due to some sparse autoencoder features only existing in some layers, we could only consider intervention topics { ‘beauty’, ‘coffee’, ‘dogs’}. We hold the hyperparameter $\alpha$ that controls for intervention “strength” constant across all layers. Note that this is NOT equivalent to holding the normalized edit distance constant, as shown in the rightmost plot.
We find that layer depth seems to have minimal effect for SAEs and probing, with the exception of the first and last layers. For steering vectors, we observe a modest increase in intervention success rate with increased layer depth and a much more drastic increase in the success rate at later layers for Logit Lens and Tuned Lens. However, as we increase $\alpha$ significantly, we find that the curves for all three methods on intervention rate shift left until the pass rate is approximately 1 at all layers. Intuitively, this makes sense, as any edits to the residual stream at layer 0 will affect the residual stream at later layers. We note that these results highlight the need to tune the intervention strength for each method, each model, and each layer - limiting their ease of use.
Figure 10: layers
A.7 Additional Example Outputs
We present additional examples of the output text for all intervention methods in Figures 11 and 12 for qualitative evaluation of intervention on the feature ‘coffee’ and ‘San Francisco’. Examples for the “Optimal intervention strength” (left column) were randomly chosen from the outputs where intervention succeeded and coherence was still relatively high. Examples for “Excessive Intervention” were randomly chosen from the outputs of the highest intervention strength tested (right column). Please see Section 4.4 for more.
Figure 11: Example outputs with intervention on “coffee” feature.
Figure 12: Example outputs with intervention on “San Francisco” feature.

### v2

@@
-Towards Unifying Interpretability and Control: Evaluation via Intervention
+
 Abstract
-With the growing complexity and capability of large language models (LLMs), a need to understand model reasoning has emerged, often motivated by an underlying goal of controlling and aligning models.
-While numerous interpretability and steering methods have been proposed as solutions, they are typically designed either for understanding or for control, seldom addressing both, with the connection between interpretation and control more broadly remaining tenuous.
-Additionally, the lack of standardized applications, motivations, and evaluation metrics makes it difficult to assess these methods’ practical utility and efficacy.
-To address the aforementioned issues, we propose intervention as a fundamental goal of interpretability and introduce success criteria to evaluate how well methods are able to control model behavior through interventions.
-We unify and extend four popular interpretability methods—sparse autoencoders, logit lens, tuned lens, and probing—into an abstract encoder-decoder framework.
-This framework maps intermediate latent representations to human-interpretable feature spaces, enabling interventions on these interpretable features, which can then be mapped back to latent representations to control model outputs.
-We introduce two new evaluation metrics: intervention success rate and the coherence-intervention tradeoff, designed to measure the accuracy of explanations and their utility in controlling model behavior.
-Our findings reveal that (1) although current methods allow for intervention, they are inconsistent across various models and features, (2) lens-based methods outperform others in achieving simple, concrete interventions, and (3) interventions often compromise model performance and coherence, underperforming simpler alternatives, such as prompting, for steering model behavior and highlighting a critical shortcoming of current interpretability approaches in real-world applications requiring control.
-Code is made available for replicability.
+With the growing complexity and capability of large language models, a need to understand model reasoning has emerged, often motivated by an underlying goal of controlling and aligning models.
+While numerous interpretability and steering methods have been proposed as solutions, they are typically designed either for understanding or for control, seldom addressing both.
+Additionally, the lack of standardized applications, motivations, and evaluation metrics makes it difficult to assess methods’ practical utility and efficacy.
+To address the aforementioned issues, we argue that intervention is a fundamental goal of interpretability and introduce success criteria to evaluate how well methods can control model behavior through interventions.
+To evaluate existing methods for this ability, we unify and extend four popular interpretability methods—sparse autoencoders, logit lens, tuned lens, and probing—into an abstract encoder-decoder framework, enabling interventions on interpretable features that can be mapped back to latent representations to control model outputs.
+We introduce two new evaluation metrics: intervention success rate and coherence-intervention tradeoff, designed to measure the accuracy of explanations and their utility in controlling model behavior.
+Our findings reveal that (1) while current methods allow for intervention, their effectiveness is inconsistent across features and models, (2) lens-based methods outperform SAEs and probes in achieving simple, concrete interventions, and (3) mechanistic interventions often compromise model coherence, underperforming simpler alternatives, such as prompting, and highlighting a critical shortcoming of current interpretability approaches in applications requiring control.
 1 Introduction
@@
 Many interpretability methods aim to address this problem by analyzing model representations, attempting to understand their underlying computational and reasoning processes in order to ultimately control model behaviour.
-While many of these methods, and interpretability as a field more broadly, claim control and intervention as abstract goals and present compelling qualitative results demonstrating that intervention is possible (for example, Anthropic’s Golden Gate Claude Anthropic ; Templeton et al.
-(2024)), the link between interpretation and intervention is often tenuous in practice, as many methods are not explicitly tailored for both.
-Furthermore, even fewer are thoroughly and systematically evaluated for the ability to control model outputs beyond simple qualitative examples.
+While many of these methods, and interpretability as a field more broadly, claim control and intervention as abstract goals and present compelling qualitative results demonstrating that intervention may be possible in certain cases (for example, Anthropic’s Golden Gate Claude Anthropic ; Templeton et al.
+(2024)), the link between interpretation and intervention is tenuous in practice, and many methods are not explicitly tailored for both.
+Furthermore, even fewer are thoroughly and systematically evaluated for the ability to control model outputs beyond qualitative examples.
 We believe the reason for this is threefold.
-First, interpretability methods produce explanations in disparate feature spaces, such as the token vocabulary, probe predictions, or learned auto-interpreted features, hindering comparisons across methods.
+First, interpretability methods produce explanations in disparate feature spaces, such as token vocabulary, probe predictions, or learned auto-interpreted features, hindering comparisons across methods.
 Second, there exists a “predict/control discrepancy” (Wattenberg & Viégas, 2024), where the features identified by interpretability methods for predicting behavior are not the same as those used for steering it.
@@
 Figure 1: Our proposed intervention framework, which encodes model latent representations, $x$, into human-interpretable features, $z=xD$, that can then be perturbed to $z^{\prime}$ and decoded back into counterfactual latent representations, $\hat{x}^{\prime}$.
-In this work, we view intervention as a fundamental goal of interpretability, and propose to measure the correctness of interpretability methods by their ability to successfully edit model behaviour.
-In particular, we focus on sparse autoencoders  (Cunningham et al., 2023; Rajamanoharan et al., 2024a; Templeton et al., 2024; Bricken et al., 2023; Gao et al., 2024), Logit Lens  (nostalgebraist, ; Dar et al., 2023), Tuned Lens  (Belrose et al., 2023; Din et al., 2023), and linear probing  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022), and benchmark them with steering vectors and prompting as baselines for intervention.
+In this work, we view intervention as a fundamental goal of interpretability, and propose to measure both the correctness and the utility of interpretability methods by their ability to successfully edit model behaviour.
+In particular, we focus on sparse autoencoders  (Cunningham et al., 2023), Logit Lens  (nostalgebraist, ; Dar et al., 2023), Tuned Lens  (Cunningham et al., 2023; Rajamanoharan et al., 2024; Templeton et al., 2024; Bricken et al., 2023; Gao et al., 2024), and linear probing  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022), and benchmark them with steering vectors and prompting as baselines for intervention.
 In order to enable comparison across these various methods, we first unify and extend the methods as instances of an abstract encoder-decoder framework, where each method encodes uninterpretable latent representations of language models into human-interpretable features and the decoder of the framework inverts this mapping, allowing us to reconstruct a latent representation from the features.
@@
 The unifying feature interpretation and intervention framework allows us to propose two standard metrics for evaluating mechanistic interpretability methods: (1) intervention success rate, which measures how well intervening on an interpretable feature causally results in the desired behavior in the model outputs, and (2) coherence-intervention tradeoff, which measures how well the causal interventions succeed without damaging the coherence of the model’s outputs.
-We evaluate Logit Lens, Tuned Lens, sparse autoencoders, and linear probes, for these metrics on GPT2-small, Gemma2-2b, and Llama2-7b, comparing them to simpler but uninterpretable baselines of steering vectors and prompting.
-Our results show that while existing methods allow for intervention, they are inconsistent across features and models.
-Furthermore, lens-based methods outperform all other methods including sparse autoencoders for simple, concrete features.
-This is likely due to frequent mislabeling of the learned SAE features, where a feature labeled “references to coffee” actually seems to encode for ‘beans’ and ‘coffee beans’ more than ‘coffee’.
-Similarly, probes and steering vectors are prone to learning spurious correlations in their training data, whereas Logit Lens and Tuned lens do not suffer from the same sensitivities.
-Finally, we show that intervention often comes at the cost of model output coherence, underperforming simple prompting baselines, presenting a critical shortcoming of existing methods in real-world applications that require control and intervention.
-We conclude this work with detailed takeaways about the strengths and weaknesses of each method, including discussion of which methods are optimal for specific intervention topics, which are best to use out of the box, and which hold the most promise for future development.
+We evaluate Logit Lens, Tuned Lens, sparse autoencoders, and linear probes for these metrics on GPT2-small, Gemma2-2b, Llama2-7b, and Llama3-8b, comparing them to simpler but uninterpretable baselines of steering vectors and prompting.
+Our results show that while existing methods allow for intervention, their effectiveness is inconsistent across features and models.
+Furthermore, lens-based methods outperform all other methods, including sparse autoencoders, for simple, concrete features, likely due to the spurious correlation learned by probes and steering vectors and the high error rate in SAE feature labeling pipelines.
+We further show that intervention often comes at the cost of model output coherence, underperforming simple prompting baselines, presenting a critical shortcoming of existing methods in real-world applications that require control and intervention.
+We conclude this work with some case studies of intervention on complex and safety-relevant features, along with detailed takeaways about the strengths and weaknesses of each method, including discussion of which methods are optimal for specific intervention topics, which are best to use out of the box, and which hold the most promise for future development.
 Our main contributions include:
-  • In Section 3.1, we present an encoder-decoder framework that unifies four popular mechanistic interpretability methods: sparse autoencoders, logit lens, tuned lens, and probing.
-  • In Section 3.2, we propose two evaluation metrics for encoder-decoder interpretability methods, namely (1) intervention success rate; and (2) the coherence-intervention tradeoff to evaluate the ability of interpretability methods to control model behavior.
-We also provide a dataset for benchmarking state-of-the-art LLMs on these metrics.
-  • In Section 4, we perform experimental analysis on GPT-2, Gemma2-2b, and LLama2-7b, and present detailed takeaways comparing interpretability- and control-based methods.
-We hope that this paper encourages future work in the interpretability community to make explicit the intent and application of methods, particularly in reference to control and intervention, and to evaluate and benchmark causality and utility in application.
+  • In Section 3.1, we present a unifying framework for four popular interpretability methods: sparse autoencoders, logit lens, tuned lens, and probing.
+To faciliate this, we extend logit lens and tuned lens methods with decoders to allow for intervention.
+  • In Section 3.2, we propose two evaluation metrics for encoder-decoder interpretability methods, namely (1) intervention success rate and (2) the coherence-intervention tradeoff to evaluate the ability of interpretability methods to control model behavior, and design an open-ended prompt dataset for benchmarking interpretability methods.
+  • In Section 4, we perform experimental analysis on GPT-2, Gemma2-2b, Llama2-7b, and Llama3-8b, and present detailed takeaways comparing interpretability- and control-based methods.
+Overall, this paper takes a key step in establishing systematic benchmarks for mechanistic interpretability methods, making progress towards a previously stated open problem for the field Mueller et al.
+(2024).
+[^1] Code and data are hosted at https://github.com/AI4LIFE-GROUP/interp_interv.git.
 2 Related Work
@@
 As such, we focus primarily on methods for inspecting hidden representations, of which probes are the most commonly used  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022).
-Other methods such as nostalgebraist ; Dar et al.
+Other methods such as Logit lens nostalgebraist ; Dar et al.
 (2023) project intermediate representation into the token vocabulary space, with Belrose et al.
@@
 (2024a) unifies most of these methods into an abstracted framework for inspecting model computation.
-More recently, sparse autoencoders and dictionary learning have been explored as a solution to the uninterpretability of model neurons, particularly due to issues with polysemanticity and superposition (Elhage et al., ; Bricken et al., 2023; Cunningham et al., 2023; Bhalla et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024a; Templeton et al., 2024; Karvonen et al., 2024; Dunefsky et al., 2024).
+More recently, sparse autoencoders and dictionary learning have been explored as a solution to the uninterpretability of model neurons, particularly due to issues with polysemanticity and superposition (Elhage et al., ; Bricken et al., 2023; Cunningham et al., 2023; Bhalla et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024; Templeton et al., 2024; Karvonen et al., 2024; Dunefsky et al., 2024).
 Evaluation.
-Due to the recency of the field, standard evaluation metrics across interpretability methods have not yet been established, and similar to the broader interpretability field, evaluation is frequently ad-hoc and primarily qualitative in nature.
+Due to the recency of the field, standard evaluation metrics across interpretability methods have not yet been established, and similar to the broader interpretability field, evaluation is frequently ad-hoc and primarily qualitative in nature, with recent works pointing to the need for more causal evaluation Mueller et al.
+(2024); Saphra & Wiegreffe (2024).
 With regards to quantitative metrics, in Gao et al.
@@
 (2024); Makelov et al.
-(2024), sparse autoencoders are evaluated for reconstruction error, recovery of supervised or known features, activation precision, and the effects of ablation; however, none of these metrics measure the human interpretability or correctness of explanations.
-Sparse autoencoder evaluation has also sometimes included ratings by humans to determine the interpretability of the learned features for labeling, as done in Rajamanoharan et al.
-(2024b), but such measures do not necessarily reflect downstream utility in model control.
+(2024), sparse autoencoders are evaluated for reconstruction error, recovery of supervised or known features, activation precision, and the effects of ablation; however, none of these metrics measure the correctness of explanations or usefulness for control.
+Independent of our work, Wu et al.
+(2025) also propose a benchmark for steering methods, AxBench, to assess whether steering is a viable alternative to existing model control techniques, finding similar results to ours.
+Different from them, we consider additional lens-based interpretability methods and explore the extent to which intervention is possible without output degradation, for both simple and safety-relevant interventions.
 Causal Intervention.
-Previous literature on probing frequently evaluates learned probes and features through intervention to ensure causality and correctness, as done in Li et al.
+Previous literature on probing frequently evaluates learned probes and features through intervention to ensure causality and correctness, as done by Li et al.
 (2022); Chen et al.
@@
 (2024b) and should ideally produce the same effect but with the added claim of interpretability.
-Beyond probing, only Belrose et al.
+Geiger et al.
+(2024) unify many interpretability methods and steering through causal abstraction but do not extend or evaluate these methods for control.
+Mueller et al.
+(2024); Belrose et al.
 (2023); Chan et al.
 (2022); Olah et al.
-(2020); Templeton et al.
-(2024) consider causal intervention as a tool for evaluation.
-However, Belrose et al.
-(2023); Olah et al.
-(2020); Chan et al.
-(2022) mainly focus on mean ablation for intervention, with no exploration of the coherence of the counterfactual outputs or their utility in application.
+(2020) consider causal intervention as a tool for assessing explanation faithfulness; however, these works often do not compare between methods and do not consider intervention as a means for control, providing no exploration of the quality of the intervened outputs or their utility in application.
 Templeton et al.
-(2024) on the other hand only provides a qualitative evaluation of intervention via their ‘Golden Gate Claude’ [Anthropic ] and does not compare to other interpretability methods.
+(2024) on the other hand provides a qualitative demonstration of intervention via their ‘Golden Gate Claude’ but do not systematically measure or compare against other interpretability methods.
+Different from these works, our work aims to adapt and evaluate existing methods (notably, lens-based methods and SAEs) originally proposed as model inspection tools, for intervention.
 3 Method
@@
 $z^{\prime}$ $=\text{Edit}(z),~{}~{}\hat{x^{\prime}}=g(z^{\prime})$
-where each $z_{i}$ is a feature activation, each $i$ in $D$ corresponds to a human-interpretable feature [^1] While “features” are not necessarily well-defined in the interpretability community, we find that the second definition in Elhage et al.
-most closely aligns with our own notion of features as human-interpretable concepts that are relevant for manipulation and editing., and $\sigma$ is an activation function that is frequently the identity.
-In the case of sparse autoencoders, $D$ is a learned, overcomplete dictionary, with $N$ in range 16k - 65k for small models (up to 16M for large models), and $\sigma$ is a ReLU, JumpReLU, or ReLU + top-k activation function.
-Given that SAE features are learned, they are not immediately interpretable and must be labelled by humans or strong LLMs.
+where each $z_{i}$ is a feature activation, each $i$ in $D$ corresponds to a human-interpretable feature, and $\sigma$ is an activation function that is frequently the identity.
+In the case of sparse autoencoders, $D$ is a learned, overcomplete dictionary, with 16k - 65k features for small models (up to 16M for large models), and $\sigma$ is a ReLU, JumpReLU, or ReLU + top-k activation function.
+Given that SAE features are learned, they are not immediately interpretable and must be labelled by humans or strong LLMs after training.
 For Logit Lens, $D$ is simply the language model’s unembedding matrix, meaning each feature corresponds to a single token in the vocabulary.
 For Tuned Lens, $D$ is the exact same as Logit Lens but with a learned linear transformation applied.
-Linear probes can be thought of as a learned dictionary with $N=1$ where $\sigma$ is a sigmoid or softmax activation.
-Of all these methods, Logit Lens is the only method that does not require any training data, and sparse autoencoders are the only method that does not produce immediately interpretable features.
-Please see Figure 1 for a visual summary of these methods.
+Linear probes can be thought of as a learned dictionary with $N=1$ where $\sigma$ is a sigmoid or softmax activation and the data is labelled.
+Supervised dictionary learning essentially learns $N>>1$ probes with a loss that either enables latent reconstruction or language modeling, as in ReFT-r1 Wu et al.
+(2025).
+Of all these methods, Logit Lens is the only method that does not require any training data, and sparse autoencoders are the only method that do not produce immediately interpretable features.
+For a visual summary of this framework, see Figure 1.
 Interpretable features to counterfactual latent vectors.
-While producing explanations is straightforward for each of these methods, intervening on model representations using the information provided by explanations is not as simple.
+While producing explanations is straightforward for each method, intervening on model representations using the information provided by explanations is not as simple.
 Doing so requires defining a reverse mapping from the explanations to the latent representations of the model, which is only explicitly done by sparse autoencoders.
-We defined the inverse mappings for the rest of the functions as follows.
+We extend lens-based methods and probing by defining inverse mappings for them as follows.
 To map Logit Lens’s explanations back into the model’s latent space, we would ideally apply the inverse of the unembedding matrix to $z$; however, in practice this is often ill-conditioned due to the dimensionality of $D$.
@@
 Notably, both of these methods only require a simple linear transformation to go back-and-forth between latent vectors and explanations.
-For probing, $D^{-1}$ can be defined such that $\hat{x}^{\prime}=x+\theta$, as proposed by Chen et al.
-(2024) which adds the linear probe’s weights to the latent representation, bypassing the need for optimization.
+For probing, an inverse mapping $D^{-1}$ is not strictly necessary, as all interventions can be performed directly on $x$ instead of $z$, as done by Chen et al.
+(2024); however, an inverse mapping can be designed to maximally recover $x$ from $z$, as shown in Figure 1.
 Sparse autoencoders have a well-defined backwards mapping through the SAE decoder, which is frequently linear in practice and often the transpose of the encoder weights.
-Thus, we are now able to decode $z$ into $\hat{x}$ for all methods.
 Intervening on interpretable features.
@@
 For probing and steering vectors, $\hat{x}^{\prime}=x+\alpha*v$, where $v$ is the steering vector or the weights of the linear probe.
-Note that $\alpha$ is a hyperparameter that must be tuned for each method, model, and sometimes even intervention feature and thus cannot be used to compare the effects of interventions across methods.
-In order to do so, we can instead measure the normalized difference between the latent vectors $x$ and $\hat{x}^{\prime}$, to better understand the relationship between the degree of intervention performed and our various evaluation metrics.
-We also note that $\hat{x}$ and $\hat{x}^{\prime}$ are not necessarily in-distribution for the language model, but due to the additive nature of the residual stream, the linear representation hypothesis, and the linear nature of the decoder/unembedding matrix, we believe that such interventions may still be principled in practice (see Park et al.
+Note that $\alpha$ is a hyperparameter that must be tuned for each method and model, and thus cannot be used to compare the effects of interventions across methods.
+In order to do so, we can instead measure the normalized difference between the latent vectors $x$ and $\hat{x}^{\prime}$, to characterize the strength of the intervention.
+We also note that $\hat{x}$ and $\hat{x}^{\prime}$ are not necessarily in-distribution for the language model, but due to the additive nature of the residual stream and the linear representation hypothesis, we believe that such interventions may still be principled in practice (see Park et al.
 (2023) for more on the linear representation hypothesis and intervention).
-Open-ended generation.
-In order to generate open-ended text after intervening on the explanation, we edit the corresponding representations in place, as is common practice with prior steering methods.
-Formally, the representation $x_{t}$ at token position $t$ and layer $l$ is edited to be $\hat{x_{t}}^{\prime}$, ensuring a causal effect on all ensuing tokens $x_{t+1},x_{t+2},...,x_{T}$.
-3.2 Comparative Evaluation Across Methods and Models
-Given the overall lack of standardized evaluation of mechanistic interpretability methods, we intend for this work to serve as a starting point for systematic evaluation by evaluating methods in simple, easy-to-measure contexts.
+3.2 Evaluation Across Methods and Models
+Given the overall lack of standardized evaluation of mechanistic interpretability methods, we intend for this work to serve as a starting point for systematic evaluation by testing methods in simple, easy-to-measure contexts.
 In particular, we think of our evaluations as measuring a kind of upper bound for these methods: in the easiest of settings, how well do existing methods work?
-Explanation Correctness
+Explanation Correctness.
 We first propose metrics to evaluate the correctness of explanations and interventions.
-More specifically, to test whether a single feature of an explanation $z_{i}$ is correct, we intervene on that feature to produce $z_{i}^{\prime}$ and decode $z^{\prime}$ to $\hat{x}^{\prime}$.
-We can then hypothesize about the corresponding change in the model’s output when generating from $\hat{x}^{\prime}$ beforehand, as it should match the intervention made to produce $z^{\prime}$.
+More specifically, to test whether a single feature of an explanation $z_{i}$ is correct, we intervene on that feature to produce $z_{i}^{\prime}$ and decode $z^{\prime}$ to $\hat{x}^{\prime}$, which should generate text that matches the intervention made to produce $z^{\prime}$.
 For example, if feature $i$ encodes the concept “references to Paris,” increasing the value of $z_{i}$ should result in increases to references of Paris in the model’s output.
-From this, we propose a metric of Intervention Success Rate, which measures if increasing the activation of feature $i$ by increasing $z_{i}$ results in the appropriate increase of the feature $i$ in the model’s output.
+From this, we propose a metric of Intervention Success Rate, which measures if increasing activation $z_{i}$ results in the appropriate increase of the feature $i$ in the model’s output.
 To evaluate a continuous relaxation of this, we can also similarly measure the probability assigned to tokens relating to feature $i$.
@@
 Importantly, both of these metrics can be thought of as measuring the causal fidelity of the features highlighted by explanations.
-Usefulness of Intervention Methods
-[…]
-An Open-ended Evaluation Dataset.
+Figure 2: Evaluation of the Intervention Success Rate with respect to edit distance for each method on four models for the simple intervention topics.
+Note that normalized edit distance is a proxy for intervention strength that is comparable across methods.
+Logit Lens generally outperforms all other methods.
+Usefulness of Intervention Methods.
+While intervention is a useful method for evaluating the correctness of explanations, it is also a desideratum of its own and a frequent motivation for many explanation methods.
+For example, methods are often developed for the purpose of de-biasing model outputs or increasing model safety, either by localizing bad behavior or identifying it at inference time, thus allowing for targeted edits to be made.
+However, a lack of this direct connection between interpretation and model intervention has led to illusory results in prior literature (Hase et al., 2024; Wattenberg & Viégas, 2024).
+By directly and explicitly measuring how effective interpretability methods are at allowing for targeted intervention or steering, we can avoid such failure cases.
+Importantly, intervention is only useful if the language model retains its overall performance and still satisfies the purpose of the query as well as the intervention.
+Thus, we want to evaluate whether interpretability methods can steer model outputs towards feature $i$ without damaging the model.
+We define Coherence as the grammatical correctness, consistency, and relevance to the prompt of the generated text, which can be measured by querying an appropriate oracle, such as a human or strong LLM.
+Similarly, we can also measure the Perplexity of the intervened outputs with respect to a strong language model.
+In practice, we use Llama3.1-8b for both of these metrics, as it is reasonable sized, high-performing, and open source, allowing for the measurement of perplexity.
+We compare coherence scores given by Llama3.1-8b to those generated by human raters as well as a rules-based grammar checker to ensure efficacy of our LLM-as-a-judge setup in Table 1.
 […]
 4 Experiments
-In this section, we evaluate the five methods on Gemma2-2b Team et al.
-(2024), Llama2-7b Touvron et al.
-(2023), and GPT2-small Radford et al.
-(2019).
-We first measure the completeness of the explanations produced in Section 4.2.
-We then benchmark the methods on the metrics described in Section 3.2, providing quantitative results in Sections 4.3 and 4.4 and qualitative results in Section 5.
-Finally, we present an analysis of the empirical directional similarities between methods in Section 4.5, and analyses of the efficacy of methods at different model depths in Appendix A.6.
+Figure 3: Intervened output coherence measured with respect to intervention success rate.
+The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.
+In this section, we evaluate the four interpretability methods on our metrics from Section 3.2.
+We also provide case studies of intervention on more complex and safety-relevant concepts in Section 4.4.
+Finally, we present an analysis of the empirical alignment between methods in Section 4.5 and a comparison of the strengths and weaknesses of each method in Section 4.6.
+Additional experiments relating to latent reconstruction error and intervention efficacy across model layers are in Appendix A.1 and A.4.
 4.1 Implementation Details
 Intervention Topics.
-We pick 10 intervention topics that all relate to references to specific words or phrases: {‘beauty’, ‘chess’, ‘coffee’, ‘dogs’, ‘football’, ‘New York’, ‘pink’, ‘San Francisco’, ‘snow’, ‘yoga’}.
-These simple, low-level features are ideal for evaluation through intervention for four key reasons: first, measuring the presence of a word or phrase is much easier than measuring a high-level abstract concept such as sycophancy, second, these features were available across many layers for the Gemma2-2b and GPT2-small sparse autoencoders, third, the features necessarily exist in the Logit Lens unembedding dictionary, and finally, datasets that are labelled for the presence of these features are very straightforward to collect for generating steering vectors and probes.
-As such, we can easily compare intervention on these features across methods and also easily measure intervention success by checking if the given word/phrase exists in the model’s output.
+We choose 10 intervention topics that all relate to references to specific words or phrases: {‘beauty,’ ‘chess,’ ‘coffee,’ ‘dogs,’ ‘football,’ ‘New York,’ ‘pink,’ ‘San Francisco,’ ‘snow,’ ‘yoga’}, generalizing ‘Golden gate Claude’-style interventions.
+These simple, low-level features are ideal for evaluation through intervention for four key reasons: first, measuring the presence of a word or phrase is much easier than measuring a high-level abstract concept such as sycophancy, second, these features were present in the pretrained and labelled sparse autoencoders we studied, third, the features necessarily exist in the Logit Lens unembedding dictionary, and finally, datasets that are labelled for the presence of these features are very straightforward to collect for generating steering vectors and probes.
+As such, we can easily compare interventions on these features across all interpretability methods and measure intervention success by checking if the given word/phrase exists in the model’s output.
+Steering vectors and probing.
+We implement steering vectors with Contrastive Activation Addition (CAA) Rimsky et al.
+(2023) with a few simple modifications.
+Where in CAA, the difference between contrastive pairs is taken only at the last token, we find that averaging across the token dimension and taking the difference between those averages yields much better results.
+This is due to the fact that in CAA, the only difference between representations occurred in the token position of the answer letter, or the last token; however, in our case the information related to the intervention feature could be present at any token.
+Example contrastive data pairs were hand-generated by the authors and then used to prompt ChatGPT to create a total of 200 pairs of sentences.
+All data was verified by the authors and is made available in the accompanying codebase.
+These contrastive pairs were also used to train the linear probes, using the implementation from Chen et al.
+(2024).
+All probes reached train and test accuracies of 100% across all models and intervention topics.
+Sparse autoencoders and supervised dictionaries.
+We focus specifically on sparse autoencoders trained to interpret the residual stream of transformer models.
+We use the SAELens library from Bloom (2024) for GPT2-small and Llama3-8b and the Gemma Scope SAEs Lieberum et al.
+(2024) for Gemma2-2b.
+SAE feature labels were found via Neuronpedia Lin & Bloom (2023), which allows users to search through fully trained SAEs and their auto-interpretation labelled features.
+We also evaluate the Rank-1 Representation Finetuning (ReFT-r1) supervised dictionaries released by Wu et al.
+(2025), which have features that directly correspond to the SAE features for Gemma2-2b.
+Note that dictionaries were only released for layer 20 of Gemma2-2b, so we cannot present evaluation for other layers or models.
+4.2 Intervention Success Across Models
+As described in Section 3.2, in order to evaluate the correctness of explanations, we measure the causal effects of intervening on specific features of each explanation.
+For a given feature or intervention topic $i$, we see if increasing the activation of that feature results in an increase of the feature in the model’s output for the ten simple intervention topics.
+In order to compare across methods, which all have different explanation feature spaces and scales, we measure the success of interventions as a function of the norm of the distance between the edited latent representation $\hat{x}^{\prime}=g(Edit(f(x)))$ and the original latent representation $x$: $||\hat{x}^{\prime}-x||/||x||$.
+Results for intervention success rate are shown in Figure 2 and results for intervened token probability can be found in Appendix A.5.
+Across methods and models, we find that by increasing intervention strength, or the magnitude of the edit to the latent representation, intervention success rate first improves and then levels out, as expected.
+However, we unexpectedly find that Logit lens and Tuned lens generally have the highest intervention success rate, regardless of the normalized edit distance, except when compared to ReFT-r1 on Gemma2-2b.
+Furthermore, we find that SAEs, probes, and steering vectors require significantly larger edits in order to achieve reasonable intervention success.
+Note that the minimal edit distance for SAEs is nonzero, as SAE reconstruction incurs a significant error, as explored in Appendix A.1.
+In general, we believe the lower performance of SAEs is due to heavy noise in the labels of features.
+For example, a feature labelled ‘references to coffee’, is sometimes actually a feature that encodes for references to ‘beans’ and ‘coffee beans’, and thus only sometimes increases mentions of ‘coffee’.
+Probes and steering vectors also have suboptimal performance, often due to learning of spurious correlations in the training data rather than the true intervention feature.
+4.3 Effects of Intervention on Output Quality
+We next measure the coherence of the intervened output text produced by each method to ensure that intervention through interpretability methods is possible without damaging the utility of the model.
+We measure coherence as described in Section 3 as a function of the intervention success rate in Figure 3 to characterize the tradeoff between intervention success and output coherence.
+Results for coherence as a function of normalized latent edit distance, $||\hat{x}^{\prime}-x||/||x||$, are in Appendix A.3.
+We visualize the mean of coherence scores for the clean model outputs with solid black horizontal lines, the same as those shown in Figure 7, with a buffer of ±$\pm 1$ around the mean in dashed lines.
+We also consider a prompting baseline, where we simply prompt the language model to talk about the intervention topic, to better understand the optimal coherence possible while satisfying the intervention.
+This is shown by the teal stars in Figure3.
+Prompting was infeasible for GPT2-small as it was not instruction tuned.
+Also, note that the intervention success rate approaches 100% with prompting as the number of generated tokens increases; however, seeing as we only generate 30 tokens, the success rate may be lower than expected.
+Our experiments reveal that while interpretability methods may seem to provide reasonable tradeoffs between intervention success and coherence at first glance, they all underperform the simplest baseline of just prompting the model.
+Furthermore, Logit lens and Tuned lens significantly outperform all other methods when intervening on these simple topics, with intervention success rates of around 0.5 and 0.6 respectively for outputs within one point of deviation from the mean coherence score of the clean model.
+All other methods exhibit much less desirable Pareto curves, regardless of model size or intervention feature.
+Table 1: Correlation between human raters (left) and an LLM rater (Llama3-8b) for coherence or a rules-based grammar checker (right). All three raters are highly correlated with one another.
+LLM Rater vs
+Human Rater
+LLM Rater vs
+Error Checker
+Pearson r $r^{2}$ Pearson r $r^{2}$
+Llama3-8b 0.94 0.75 -0.96 0.92
+Llama2-7b 0.80 0.68 -0.85 0.73
+Gemma2-2b 0.80 0.67 -0.78 0.75
+GPT2-sm 0.71 0.67 -0.86 0.74
+Verifying Coherence.
+In order to validate the coherence scores generated through our LLM-as-a-judge setup with Llama3-8b, we verify the coherence scores with human raters.
+Participants blindly rated 100 outputs for each model, and we measured the correlation between these human ratings and LLM ratings, as shown in Table 1.
+We find high consistency between both, with particularly high correlation coefficients and $r^{2}$ values for the larger models.
+We also check the validity of the coherence ratings by comparing with an alternative metric that measures the number of grammatical errors in the intervened output via a rule-based grammar checker.
+In particular, we use LanguageTool to determine the number of errors in each output, which has thousands of rules relating to grammar, typos, capitalization errors, and more.
+As expected, there is a high negative correlation between the two, indicating that outputs with more errors are less coherent.
+However, we note that the number of grammatical errors is not an ideal metric, as it does not assess whether the text generation pertains to the prompt, which an LLM rater can do.
+Qualitative Examples.
+We present examples of intervention outputs in Figure 5 for the feature ‘yoga,’ with more examples in Appendix A.7.
+We highlight outputs where intervention succeeded with minimal degradation in coherence in “Optimal Intervention Strength” (left column) as well as generations from the highest intervention strength tested (right column) “Excessive Intervention.” Note that intervention results in repetition at very high intervention strengths for all methods; however, only Logit Lens and Tuned Lens result in repetition of tokens related to ‘yoga.’
+Figure 4: Relationship between intervention success rate and coherence for three complex features: religious references (top), gendered language (middle), and French language (bottom) for Gemma2-2b (left) and Llama3-8b (right).
+Figure 5: Examples of intervened model outputs for intervention feature ‘yoga’ at both the optimal intervention strength (left) and the maximum intervention strength tested (left).
+Outputs degrade into incoherent repetition at high intervention strength for all methods.
+4.4 Intervention on Complex Features
+While the aforementioned simple features allow for rigorous evaluation across methods, in practice, users often want to control or steer much more complex concepts.
+To investigate the feasibility of interventions in more interesting and realistic settings, we present results for three more complex concepts: (1) religious speech, with direct or implicit references to a given religion, (2) gendered-language, or the ability to preferentially generate text related to a specific gender, and (3) the French language, where the model should generate text in French even when given an English prompt.
+These concepts were chosen from features known to exist in the pretrained and labeled SAEs we evaluated, which had a female gendered-language concept and a Christianity- and Islam-related religion concept.
+[^2] ReFT-r1 did not have a feature that directly corresponded to French language, so we consider the closest successful feature available: “French connective and referential pronouns.” Intervention with Logit lens was performed by selecting ten relevant tokens as the features for the gender and religion concepts, such as {‘she’, ‘her’, ‘herself’} and {‘Christ’, ‘Allah’, ‘holy’}, but intervention could not be performed for the French concept.
+Intervention success is measured with keyword detectors for the first two concepts and the NLP-based language detector package polyglot Al-Rfou (2022) for the French concept.
+As shown in Figure 4, we find that even for these more complex concepts, prompting generally outperforms interpretability-based interventions, particularly for the French language concept.
+Furthermore, no other method performs consistently across models or features, highlighting the unreliability of existing interpretability methods for controlling complex and safety-relevant concepts.
+However, we generally found that interventions made by Logit lens were most successful in incorporating the intervention into a plausible completion of the prompt.
+For example, one intervened output was “Whenever I’m outdoors and in nature, I always have a camera with me.
+I love taking pictures of God’s beautiful creation.
+I’m a big fan of Jesus and I love spending…”, where we can see that the religion concept is integrated well with the response to the prompt.
+For most other methods, such as SAEs, interventions either immediately degraded model performance by inducing severe repetition (e.g.
+“I had a conversation recently about the last and final prophet of the last and final of the last and the last and the last and”) or they were either unrelated to the prompt or simply incorrect (e.g.
+“Tell me an interesting fact about a musical instrument.
+For example, did you know that the piano is actually a Christian Muslim?”).
+4.5 Intervention Similarity Between Methods
+Figure 6: Cosine similarity between methods’ intervention directions in model latent space across methods.
+Given that these methods all result in linear edits that should correspond to the same feature, ideally their interventions should all point in the same direction in the model’s latent space.
+We evaluate the empirical similarity between methods by measuring the cosine similarity between edit directions, $\hat{x}^{\prime}-x$, for each intervention topic.
+The average cosine similarity between these vectors for each intervention topic is shown in Figure 6.
+We find that Logit Lens and Tuned Lens are highly similar, as expected.
+Similarly, steering vectors and probe weights tend to lie in similar directions, likely due to the same underlying data used to train both.
+Most interestingly, we find that sparse autoencoders tend to intervene in somewhat similar directions to steering vectors and probes and have near orthogonal directions to Logit Lens and Tuned Lens, even when interventions succeed for all methods.
+We speculate that sparse autoencoders may be more similar to probes and steering vectors because the three methods may have a bias toward representing past information and tokens, due to their training and labelling algorithms, also noted by Gur-Arieh et al.
+(2025).
+Logit lens and Tuned lens, on the other hand, are designed to reveal information about the next token specifically, given that they are early-decoding strategies and thus may contain more information about model outputs rather than inputs.
+4.6 High-Level Comparison of Methods
+We present the following characterization of the strengths and weaknesses of the interpretability and steering methods.
+Logit Lens is easy to use, requires no training, and maps features directly to vocabulary tokens, making it highly interpretable, and we find that it generally has high causal fidelity as well.
+However, its predefined, static features are limited and rudimentary, preventing its utility in many real-world or safety-critical applications.
+Tuned Lens shares these traits but requires an additional learned affine transformation, although many are already open-sourced.
+Sparse autoencoders learn a wide variety of both low- and high-level features; however, the post-hoc labeling of features is often highly erroneous and leads to lower causal fidelity in practice.
+They are also extremely data- and compute-intensive, and there is no guarantee that any given desired feature will exist in the SAE’s dictionary.
+Supervised dictionaries can contain any arbitrary feature but also require significant data and compute to learn, along with an exhaustive list of relevant features.
+Importantly, they can be trained with intervention in mind, increasing their utility for steering while maintaining interpretability.
+Probing and steering vectors also allow for feature specification but are prone to learning spurious correlations, leading to low causal fidelity, and both require labelled data.
+As such, we believe that lens-based methods are most useful for providing high-fidelity explanations but are likely not effective solutions for steering in real applications, for which steering vectors and prompting are more promising but require careful oversight and refinement to ensure efficacy due to their uninterpretability.
+Finally, supervised dictionaries seem to hold a nice balance between interpretation- and control-based methods if users are not compute- and data-bound, proving to be more promising that unsupervised methods such as SAEs.
+5 Conclusion
+While interpretability methods show great promise in understanding large language models, the correctness of their explanations is less clear.
+Do these explanations reveal truth about model computation or simply fool human researchers?
+We believe that systematic benchmarking of explanations is critical to answer this question.
+Our work makes progress towards this goal, and answers this question somewhat negatively, showing that current explanations are less accurate than expected.
+Our work also raises questions regarding the utility of such methods, as we find that prompting outperforms current interpretability methods in its ability to steer models, without requiring any training, data, or access to model weights.
+We hope future work can address these shortcomings of current methods, paving way toward interpretability methods that are faithful and provide actionable insights for improving and controlling models.
+Acknowledgements
+The authors would like to thank Fred Zhang and Lucas Dixon for their helpful discussion and comments on the manuscript.
+This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, AI2050 Early Career Fellowship by Schmidt Sciences, and faculty research awards from Google, OpenAI, Adobe, JPMorgan, Harvard Data Science Initiative, and the Digital, Data, and Design (Dˆ3) Institute at Harvard.
+UB is funded by the Kempner Institute Graduate Research Fellowship.
+The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.
+Impact Statement
+This paper aims to improve the evaluation of interpretability methods, ensuring that they are both faithful and useful in practice.
+We are particularly interested in facilitating discussions around how to make interpretability methods more reliable to improve model fairness and safety, both of which can have significant societal implications.
+Ultimately, we hope that clear evaluation of interpretability methods will ensure that methods produce desired results, decreasing unintended consequences due to spurious correlations or unfaithful explanations.
+However, we note that interpretability can just as easily be used to make models more harmful than helpful.
+Appendix A Appendix
+A.1 Additional Evaluations: Sanity Checking Explanation Reconstructions
+Before testing these methods for their ability to intervene, we first want to evaluate the completeness of the explanations and the effect of replacing $x$ with $\hat{x}$ without any intervention or editing.
+We do so by measuring the normalized latent reconstruction error: $\texttt{Error}=||\hat{x}-x||/||x||$ where $\hat{x}=g(f(x))=g(z)$.
+This error is a key part of the loss function that sparse autoencoders are trained on and measures the information loss incurred by mapping between the language model’s latent space and the interpretable feature space.
+Given that steering vectors and linear probes do not output complete explanations, we only measure this error for the other three methods, as shown in Table 2, where we see that errors vary a lot across models but most methods are relatively consistent in their error, with the exception of the GPT2-small sparse autoencoders.
 Table 2: Normalized latent reconstruction error without intervention.
@@
 SAEs 0.38 1.64
-Steering vectors and probing.
-We implement steering vectors as described in Rimsky et al.
-(2023) with a few simple modifications.
-Where in Rimsky et al.
-(2023), the difference between contrastive pairs is taken only at the last token, we find that averaging across the token dimension and taking the difference between those averages yields much better results.
-This is due to the fact that in Rimsky et al.
-(2023), the only difference between representations occurred in the token position of the answer letter, or the last token; however, in our case the information related to the intervention feature could be present at any token.
-Contrastive data pairs were generated by prompting ChatGPT for 200 pairs sentences or phrases about the desired intervention topic and about other topics within the same genre.
-For example, we used the prompts “Can you please generate 100 short sentences or phrases with the color ‘pink’ in them?” to create positive samples and “Can you also generate 100 phrases or sentences with other colors (not pink) in them?” to create negative ones.
-We also use these contrastive pairs to train the linear probes, using the implementation from Chen et al.
-(2024).
-All probes reached train and test accuracies of 100% across all models and intervention topics.
-Sparse autoencoders.
-We focus specifically on sparse autoencoders trained to interpret the residual stream of transformer models.
-We use the SAELens library from Bloom (2024b) for GPT2-small and the Gemma Scope SAEs released by Google in Lieberum et al.
-(2024) for Gemma2-2b.
-For both models, SAE feature labels were found via Neuronpedia Lin & Bloom (2023), which allows users to interact with, explore, and search for auto-interpretation labelled learned features.
-Note that the Gemma SAEs are of the JumpReLU architecture from Rajamanoharan et al.
-(2024b) whereas the GPT2-small SAEs are standard ReLU + L1 regularization SAEs, as noted in Bloom (2024a).
-4.2 Sanity Checking Explanation Reconstructions
-Before testing these methods for their ability to intervene, we first want to evaluate the completeness of the explanations and the effect of replacing $x$ with $\hat{x}$ without any intervention or editing.
-We do so by measuring the normalized latent reconstruction error: $\texttt{Error}=||\hat{x}-x||/||x||$ where $\hat{x}=g(f(x))=g(z)$.
-This error is a key part of the loss function that sparse autoencoders are trained on and measures the information loss incurred by mapping between the language model’s latent space and the interpretable feature space.
-Given that steering vectors and linear probes do not output complete explanations, we only measure this error for the other three methods, as shown in Table 1, where we see that errors vary a lot across models but most methods are relatively consistent in their error, with the exception of the GPT2-small sparse autoencoders.
-We also measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7.
+A.2 Additional Evaluations: Coherence of Method Outputs without Intervention
+We measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7, which we can compare to the baseline of the clean model outputs (labelled ‘Clean’ and shown in black).
 We find that the coherence of the outputs generated by the reconstructed latents generally matches the coherence of the clean model outputs.
 We use a deviation of ±$\pm 1$ around the mean of clean output coherence scores as a threshold for future evaluations, shown in the dashed lines.
-Figure 2: Evaluation of the Intervention Success Rate for each method.
+Figure 7: Histogram of coherence scores for clean model outputs (Clean) and for the models where $x$ is replaced by $\hat{x}$ without any intervention for Logit Lens, Tuned Lens, and SAEs.
+Dashed lines show the mean for each distribution.
+A.3 Additional Evaluation: Coherence of Intervention with respect to Edit Distance
+We measure the coherence of the intervened output text produced by each method to ensure that intervention through interpretability methods is possible without damaging the utility of the model.
+We measure coherence as described in Section 3 as a function of normalized latent edit distance, $||\hat{x}^{\prime}-x||/||x||$ in Figure 8.
+We find that even the smallest interventions made with Logit lens and Tuned lens result in significant degradation of model outputs, with a less noticeable dropoff for the other methods.
+We also plot coherence as a function of the intervention success rate in Figure 3 to characterize the tradeoff between intervention success and output coherence.
+Figure 8: Analysis of coherence of the intervened outputs, measured with Llama3.1-8b, as a measure of the edit distance or magnitude of intervention made.
+Lens-based methods suffer drastic drops in coherence with only small edits.
+A.4 Additional Evaluations: Intervention Efficacy Across Model Depth
+In order to ensure the generalizability of the above results across layer depths, we repeat all experiments for each layer of GPT2-small, as shown in Figure 9.
+However, due to some sparse autoencoder features only existing in some layers, we could only consider intervention topics { ‘beauty’, ‘coffee’, ‘dogs’}.
+We hold the hyperparameter $\alpha$ that controls for intervention “strength” constant across all layers.
+Note that this is NOT equivalent to holding the normalized edit distance constant, as shown in the rightmost plot.
+We find that layer depth seems to have minimal effect for SAEs and probing, with the exception of the first and last layers.
+For steering vectors, we observe a modest increase in intervention success rate with increased layer depth and a much more drastic increase in the success rate at later layers for Logit Lens and Tuned Lens.
+However, as we increase $\alpha$ significantly, we find that the curves for all three methods on intervention rate shift left until the pass rate is approximately 1 at all layers.
+Intuitively, this makes sense, as any edits to the residual stream at layer 0 will affect the residual stream at later layers.
+We note that these results highlight the need to tune the intervention strength for each method, each model, and each layer - limiting their ease of use.
+Figure 9: Analysis of intervention pass rate (left), coherence (middle) and edit distance (right) across all layers of GPT2-sm.
+We find that intervening at later layers is significantly more effective for Logit and Tuned Lens than earlier interventions, but probes, steering vectors, and SAEs are relatively invariant to the choice of layer.
+A.5 Additional Metrics: Intervened Token Probability
+Please see Section 3.2 for more details.
+We measure the probability assigned to tokens relating to feature $i$ when intervening on feature $i$.
+As such, even if a model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of feature $i$ results in any change to the model’s output at all.
+We refer to this metric as Intervened Token Probability.
+Results for Intervened Token Probability are shown in Figure 10, where we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed.
+We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
+Figure 10: Evaluation of intervention success with respect to the probabilities of the tokens corresponding to the features intervened on for each method.
 Note that normalized edit distance is a proxy for intervention intensity that is comparable across methods.
-Logit Lens outperforms all other methods across all models.
-4.3 Intervention Success Across Models
-As described in Section 3.2, in order to evaluate the correctness of explanations, we measure the causal effects of intervening on specific features of each explanation.
-For a given feature or intervention topic $i$, we see if increasing the activation of that feature results in an increase of the feature in the model’s output in the top row of Figure 2.
-In order to compare across methods, which all have different explanation feature spaces and scales, we measure the success of interventions as a function of the norm of the distance between the edited latent representation $\hat{x}^{\prime}=g(Edit(f(x)))$ and the original latent representation $x$: $||\hat{x}^{\prime}-x||/||x||$ as done in Section 4.2.
-Results for intervened token probability are in Appendix A.4.
-We find that across all three models, Logit lens and Tuned lens have the highest intervention success rate, regardless of the normalized edit distance.
-While increasing intervention strength and editing the latent representations more for steering vectors and probes results in greater intervention success for GPT2-small and LLama2-7b, the same is not true for Gemma2-2b, for which intervening with steering vectors and probes seems to have minimal effect.
-Furthermore, we find that intervening with sparse autoencoders is moderately successful, but due to the large reconstruction error for GPT2-small, even with intervention, the edit distance is much greater than that needed for Logit lens and Tuned lens.
-In general, we believe the lower performance of Sparse Autoencoders is due to heavy noise in the labels of features.
-For example, a feature labelled ‘references to coffee’, is sometimes actually a feature that encodes for references to ‘beans’ and ‘coffee beans’, and thus only sometimes increases mentions of ‘coffee’.
-Probes and steering vectors also have suboptimal performance, often due to learning of spurious correlations in the training data rather than the true intervention feature.
-When measuring token probability rather than output presence, as explored in Appendix A.4, we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed.
-We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
-4.4 Effects of Intervention on Output Quality
-We next measure the coherence of the intervened output text produced by each method to ensure that intervention through mechanistic interpretability methods is possible in applications and does not damage the utility of the model.
-We measure coherence as described in Section 3 as a function of the normalized latent edit distance $||\hat{x}^{\prime}-x||/||x||$ and with respect to the intervention success rate in Figure 4.
-We visualize the mean of coherence scores for the clean model outputs with solid black horizontal lines, the same as those shown in Figure 7, with a buffer of ±$\pm 1$ around the mean in dashed lines.
-We also consider a prompting baseline, where we simply prompt the language model to constantly talk about the intervention topic, to better understand the optimal coherence possible while satisfying the intervention.
-This is shown by the teal stars in 4.
-Given that only Llama2-7b was instruction-tuned out of the models evaluated, prompting was less successful for Gemma2-2b and infeasible for GPT2-small.
-Also note that the intervention success rate approaches near 100% with prompting as the number of generated tokens increases; however, seeing as we only generate 30 tokens, the success rate is slightly lower than expected.
-We find that for GPT2-small and Llama2-7b, intervention with most methods results in a sharp decline in text coherence, with Logit Lens and Tuned Lens having the most dramatic fall-off.
-In fact, we find that even with a full point drop in coherence from the baseline models’ coherence (denoted with the black lines), the representations can barely be edited by any methods except for Gemma2-2b.
-While at first, this seems quite concerning, as this would prevent the real-world application of these methods to scenarios necessitating high model performance, viewing coherence as a function of intervention success rate tells a different story, as shown in Figure 4.
-In fact, we see that Logit Lens and Tuned Lens have relatively good coherence-intervention tradeoffs for GPT2-small and LLama2-7b, with success rates of around 0.5 and 0.6 respectively when considering outputs within one point of deviation from the mean coherence score of the clean model.
-All other methods exhibit much less desirable Pareto curves.
-We also note that as the models increase in size and performance, with GPT2-small being the smallest and worst language model and Llama2-7b being the largest and most performant, the Pareto curves are shifted up and to the right more for most methods, suggesting that targeted intervention is more possible as models become more capable.
-Figure 3: Intervened model output coherence measured with respect to normalized edit distance.
-The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.
-Figure 4: Intervened output coherence measured with respect to intervention success rate.
-The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.
-Qualitative Examples.
-In addition to our empirical results, we present examples of the output text for all intervention methods in Figure 5 for qualitative evaluation of intervention on the feature ‘yoga’.
+A.6 Additional Metrics: Perplexity
+As described in Section 3.2, we evaluate the perplexity of the intervened generated text to measure the utility of interpretability methods for targeted intervention in 11.
+We measure this perplexity with respect to a stronger language model than the one studied, in this case with Llama3.1-8b.
+We find that the results for perplexity are generally unintuitive and do not align with the results for coherence.
+We hypothesize that perplexity is not a useful measure when text is extremely out-of-distribution with respect to normal text, and in particular when the text is highly repetitive.
+For example, if the same token is repeated 20 times, we (and other language models) might assume that the next 20 tokens would also be the same, resulting in a low perplexity even if the quality of the text is poor.
+As such, we do not consider these results to be particularly meaningful or significant.
+Figure 11: Analysis of perplexity of the intervened outputs, measured with Llama3.1-8b, as an alternative metric to Coherence.
+We find that perplexity does not align with Coherence, as highly repetitive sequences may have low perplexity despite being incoherent answers to prompts.
+A.7 Additional Example Outputs
+We present additional examples of the output text for all intervention methods in Figures 12 and 13 for qualitative evaluation of intervention on the feature ‘coffee’ and ‘San Francisco’.
 Examples for the “Optimal intervention strength” (left column) were randomly chosen from the outputs where intervention succeeded and coherence was still relatively high.
 Examples for “Excessive Intervention” were randomly chosen from the outputs of the highest intervention strength tested (right column).
-Note that intervention results in infinite repetition at very high intervention strengths for all methods; however, only Logit Lens and Tuned Lens result in repetition of tokens related to ‘yoga’.
-More examples are provided in Appendix A.7.
-Figure 5: Examples of intervened model outputs for intervention feature ‘yoga’ at both the optimal intervention strength (left) and the maximum intervention strength tested (left).
-Outputs degrade into incoherent repetition at high intervention strength for all methods.
-4.5 Intervention Similarity Between Methods
-We finally evaluate the empirical similarity between each of these methods.
-By converting each intervention into one in the latent space of the model, we can measure the cosine similarity between edit directions for each method.
-More specifically, for each method and intervention topic, we compute the expectation of $\hat{x}^{\prime}-x$ over all samples.
-This difference vector can be thought of as the direction of the intervention topic in the model’s latent space.
-We compute the cosine similarity between these vectors for each intervention topic and display the average across topics in Figure 6.
-We find that Logit Lens and Tuned Lens are highly similar, as expected.
-Similarly, steering vectors and probe weights tend to lie in similar directions, likely due to the same underlying data used to train both.
-Most interestingly, we find that sparse autoencoders tend to intervene in somewhat similar directions to steering vectors and probes.
-Even more surprisingly, they have near orthogonal directions to Logit Lens and Tuned Lens, even when interventions succeed for all methods.
-This is likely due to the high-dimensionality of model latent spaces, allowing for many vectors to be near-orthogonal.
-We speculate that sparse autoencoders may be more similar to probes and steering vectors because the three methods may have a bias toward representing past information and tokens, due to their training and labelling algorithms.
-Logit Lens and Tuned lens, on the other hand, are designed to reveal information about the next token specifically, given that they are early-decoding strategies and thus may contain more information about model outputs rather than inputs.
-Figure 6: Cosine similarity of intervention directions in model latent space across methods. Logit Lens and Tuned Lens are highly similar, as well as steering vectors and probes.
-4.6 High-Level Comparison of Methods
-Given the differences between methods, including how their features are learnt, what explanations they produce, and the causal fidelity and utility of explanations for steering, we present the following characterization of the strengths and weaknesses of each method.
-Logit Lens is easy to use, requires no training, and maps features directly to vocabulary tokens, making it highly interpretable, and we find that it generally has high causal fidelity as well.
-However, its predefined, static features are limited and rudimentary, preventing its utility for steering in many real-world or safety-critical applications where abstract features such as truthfulness and bias are more relevant.Tuned Lens shares these traits but requires an additional learned affine transformation, although many are already open-sourced.
-Sparse autoencoders learn a very large dictionary, allowing for features to be both low-level and high-level or abstract, potentially including safety-critical features, and ideally “covering” all the true underlying features learned by the model.
-Furthermore, explanations are sparse, increasing their human interpretability.
-However, the learned features are frequently not human-interpretable, leading to labeling errors and lower causal fidelity in practice.
-Furthermore, they require significant training and labeling data and compute, and there is no guarantee that a desired feature will exist in the SAE’s dictionary, also preventing their utility in safety-critical domains.
-Probing allows for feature specification, given that a labeled dataset is available, but is prone to learning spurious correlations, leading to low causal fidelity and minimal steering utility.
-Steering vectors are very similar to probes in practice, but they require no optimization and are slightly less sensitive to correlations in the training data.
-As such, we believe that lens-based methods are most useful for providing high-fidelity explanations but are likely not effective solutions for steering in real applications, for which steering vectors and prompting are more promising but require careful oversight and refinement to ensure efficacy.
-SAEs offer a balance between interpretation- and control-based methods but also suffer from a union of their problems.
-5 Conclusion
-As language models are integrated into everyday aspects of life, whether benign or life-changing, being able to understand them and ensure they behave as expected is of utmost importance.
-While interpretability methods show great promise, the correctness of their interpretations is less clear.
-In this work, we emphasize that intervention is a fundamental goal of interpretability, both to evaluate the correctness of interpretability methods as well as their utility in real-world applications.
-By unifying and extending four popular interpretability methods — sparse autoencoders, logit lens, tuned lens, and probing — into an abstracted encoder-decoder framework, we enable structured interventions on human-interpretable features, mapfping them back to model latent representations to directly control outputs.
-Through the introduction of two standardized metrics, the intervention success rate, and the coherence-intervention tradeoff, we provide a rigorous basis for evaluating the causal fidelity and utility of interpretability methods.
-Our findings reveal significant inconsistencies in intervention capability across different methods and models, with simpler lens-based approaches generally outperforming the more compute-intensive sparse autoencoders for simple, concrete interventions.
-Furthermore, non-interpretability-based approaches, such as prompting, perform best overall, demonstrating a significant gap in current interpretability approaches.
-Acknowledgments
-The authors would like to thank Fred Zhang and Lucas Dixon for their helpful discussion and comments on the manuscript.
-This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, AI2050 Early Career Fellowship by Schmidt Sciences, and faculty research awards from Google, OpenAI, Adobe, JPMorgan, Harvard Data Science Initiative, and the Digital, Data, and Design (Dˆ3) Institute at Harvard.
-UB is funded by the Kempner Institute Graduate Research Fellowship.
-The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.
-Appendix A Appendix
-A.1 Additional Implementation Details: Intervention Hyperparameter $\alpha$
+Please see Section 4.3 for more.
+Figure 12: Example outputs with intervention on “coffee” feature.
+Figure 13: Example outputs with intervention on “San Francisco” feature.
+A.8 Implementation Details: Open-ended Generation
+In order to generate open-ended text after intervening on the explanation, we edit the corresponding representations in place, as is common practice with prior steering methods.
+Formally, the representation $x_{t}$ at token position $t$ and layer $l$ is edited to be $\hat{x_{t}}^{\prime}$, ensuring a causal effect on all ensuing tokens $x_{t+1},x_{t+2},...,x_{T}$.
+A.9 Implementation Details: Intervention Hyperparameter $\alpha$
 When intervening on $z$ to get $z^{\prime}$ with Logit Lens, Tuned Lens, and SAEs, we set $z_{i}^{\prime}=\alpha*max(z)$.
@@
 We record the values of $\alpha$ used in our experiments in Table 3.
-Table 2: .
+Table 3: Values for hyperparameter $\alpha$ used to control intervention edit distance for each method and model.
 Method
@@
 Steering Vectors [2, 4, 6, 8, 10] [3, 4, 5, 6] [0.5, 3, 4, 5, 6]
-A.2 Additional Implementation Details: SAE Features
+A.10 Implementation Details: SAE Features
 As described in Section 4.1, we use the sparse autoencoders hosted on SAELens and find the relevant features with Neuronpedia’s exploration and search tools.
@@
 Many of these imperfect features still yield reasonable intervention success rates.
-Table 3: .
+Table 4: Specific SAE features used for intervention on GPT2-sm and Gemma2-2b. The feature ids and their according Neuronpedia labels are provided.
 Intervention
@@
 related cafés or establishments”
-A.3 Additional Evaluations: Coherence of Method Outputs without Intervention
-We measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7, which we can compare to the baseline of the clean model outputs (labelled ‘Clean’ and shown in black).
-We find that the coherence of the outputs generated by the reconstructed latents generally matches the coherence of the clean model outputs.
-We use a deviation of ±$\pm 1$ around the mean of clean output coherence scores as a threshold for future evaluations, shown in the dashed lines.
-Figure 7: Histogram of coherence scores for clean model outputs (Clean) and for the models where $x$ is replaced by $\hat{x}$ without any intervention for Logit Lens, Tuned Lens, and SAEs.
-Dashed lines show the mean for each distribution.
-A.4 Additional Metrics: Intervened Token Probability
-Please see Section 3.2 for more details.
-We measure the probability assigned to tokens relating to feature $i$ when intervening on feature $i$.
-As such, even if a model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of feature $i$ results in any change to the model’s output at all.
-We refer to this metric as Intervened Token Probability.
-Importantly, both of these metrics can be thought of as measuring the causality of the features highlighted by explanations.
-Results for Intervened Token Probability are shown in 8, where we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed.
-We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
-Figure 8:
-A.5 Additional Metrics: Perplexity
-As described in Section 3.2, we evaluate the perplexity of the intervened generated text to measure the utility of interpretability methods for targeted intervention in 9.
-We measure this perplexity with respect to a stronger language model than the one studied, in this case with Llama3.1-8b.
-We find that the results for perplexity are generally unintuitive and do not align with the results for coherence.
-We hypothesize that perplexity is not a useful measure when text is extremely out-of-distribution with respect to normal text, and in particular when the text is highly repetitive.
-For example, if the same token is repeated 20 times, we (and other language models) might assume that the next 20 tokens would also be the same, resulting in a low perplexity even if the quality of the text is poor.
-As such, we do not consider these results to be particularly meaningful or significant.
-Figure 9:
-A.6 Intervention Efficacy Across Model Depth
-In order to ensure the generalizability of the above results across layer depths, we repeat all experiments for each layer of GPT2-small, as shown in FIgure 10.
-However, due to some sparse autoencoder features only existing in some layers, we could only consider intervention topics { ‘beauty’, ‘coffee’, ‘dogs’}.
-We hold the hyperparameter $\alpha$ that controls for intervention “strength” constant across all layers.
-Note that this is NOT equivalent to holding the normalized edit distance constant, as shown in the rightmost plot.
-We find that layer depth seems to have minimal effect for SAEs and probing, with the exception of the first and last layers.
-For steering vectors, we observe a modest increase in intervention success rate with increased layer depth and a much more drastic increase in the success rate at later layers for Logit Lens and Tuned Lens.
-However, as we increase $\alpha$ significantly, we find that the curves for all three methods on intervention rate shift left until the pass rate is approximately 1 at all layers.
-Intuitively, this makes sense, as any edits to the residual stream at layer 0 will affect the residual stream at later layers.
-We note that these results highlight the need to tune the intervention strength for each method, each model, and each layer - limiting their ease of use.
-Figure 10: layers
-A.7 Additional Example Outputs
-We present additional examples of the output text for all intervention methods in Figures 11 and 12 for qualitative evaluation of intervention on the feature ‘coffee’ and ‘San Francisco’.
-Examples for the “Optimal intervention strength” (left column) were randomly chosen from the outputs where intervention succeeded and coherence was still relatively high.
-Examples for “Excessive Intervention” were randomly chosen from the outputs of the highest intervention strength tested (right column).
-Please see Section 4.4 for more.
-Figure 11: Example outputs with intervention on “coffee” feature.
-Figure 12: Example outputs with intervention on “San Francisco” feature.